# Anomalous dynamical Casimir effect in an expanding ring

### v2

@@ Abstract
The Casimir effect is one of the most striking macroscopic manifestations of the quantum nature of vacuum Casimir (1948). It emerges when this vacuum is subject to boundary conditions, e.g. two metallic plates separated by an electromagnetic vacuum experience an attractive Casimir force Bordag et al. (2009); Mostepanenko and Trunov (1997); Bressi et al. (2002). This effect has been tested experimentally using metals in vacuumBressi et al. (2002); Lamoreaux (1997); Mohideen and Roy (1998) and turned to be repulsive by replacing vacuum by a fluid Munday et al. (2009). The Casimir effect generalizes to the case of dielectric plates Lifshitz et al. (1992); Klimchitskaya et al. (2009); Woods et al. (2016), liquid crystals Ajdari et al. (1992), metamaterialsBoyer (1974); Kenneth et al. (2002); Henkel and Joulain (2005); Leonhardt and Philbin (2007); Rosa et al. (2008), liquids and gases Le Cunuder et al. (2018), and relativistic Zhabinskaya et al. (2008); Shytov et al. (2009); Lu (2021) and topological Grushin and Cortijo (2011); Grushin et al. (2011); Rodriguez-Lopez and Grushin (2014) matter. When the boundary conditions are time-dependent, the increase of vacuum energy and associated properties lead to the so-called dynamical Casimir effects Moore (1970); Dodonov (2020); Wilson et al. (2011); Jaskula et al. (2012); Lähteenmäki et al. (2013); Johansson et al. (2009); Vezzoli et al. (2019).
@@
The Casimir effect also affects groundstate properties in a finite geometry without boundaries. For example, when a massless field is confined in a ring of size $L$ in dimension $D=1$, the finite-size energy correction is a manifestation of the Casimir effect Francesco et al. (2012). In this paper, we consider the fate of this Casimir effect when the size of this ring is rapidly increased in time. Such a system is an analog of a 1D universe in expansion. Analogous expanding universes can be implemented experimentally for example at the edge of a quantum Hall sample with a time-dependent geometry Nambu and Hotta (2023), and in Bose-Einstein condensate by modulating either their shape Fedichev and Fischer (2004); Weinfurtner (2005); Eckel et al. (2018); Llorente and Plata (2019) or their scattering length Jain et al. (2007a).
@@
In this work we show that when the expansion of the ring is fast enough the Casimir effect has an additional contribution, distinct from the dynamical Casimir effect resulting from dynamical boundary conditions. Because of this difference, and its link to quantum anomalies Bertlmann (2000) that we now introduce, we refer to it as the anomalous dynamical Casimir effect. The onset scale of this effect is set by the curvature of spacetime which modifies the ground state properties. In the present case, this curvature of spacetime originates from the time dependence of the spatial curvature of the ring, i.e. its radius. In practice, we account for the modifications of the vacuum properties by this spacetime curvature through the conformal anomaly of the quantum field theory for the energy density and pressure, and the gravitational anomaly for the energy current and momentum Bertlmann (2000). We stress that the curved spacetime description of the dynamical Casimir effect Moore (1970); Fulling and Davies (1976) relies on a metric with a vanishing curvature Davies and Fulling (1977), unrelated to a conformal anomaly. Hence, the anomalous dynamical Casimir effect, originating from quantum field theory anomalies, is thus distinct from the dynamical Casimir effect in a quantum wire with a varying length or boundary conditions.
@@
In contrast, the anomalous dynamical Casimir effect has a similar origin to that of the Hawking or Unruh radiations. All of these phenomena originate from a curvature of spacetime Christensen and Fulling (1977); Robinson and Wilczek (2005); Banerjee and Kulkarni (2008, 2009); Gim and Kim (2015); Bermond et al. (2022), originating from either a spatial variation of the metric Unruh (1981) for the Hawking effect, or a dynamical origin for both the Unruh effect and the present anomalous dynamical Casimir effect. The anomalous dynamical Casimir effect opens an alternative to test the effect of a spacetime curvature on the vacuum properties, in addition to the experimental realization of the Hawking effect  Unruh (1981); Garay et al. (2000, 2001); Unruh and Schützhold (2007); Barcelo et al. (2001); Jain et al. (2007b); Macher and Parentani (2009); Steinhauer (2014, 2016). While the experimental signatures of these phenomena typically rely on the measurement of the particle density and its correlations Jain et al. (2007b); Steinhauer (2014, 2016); Johansson et al. (2009); Jaskula et al. (2012), here we show that a remarkable feature of the anomalous dynamical Casimir effect appears as a mechanical instability, related to the behavior of the system’s radiative pressure $p$.
@@
We consider a massless field of excitations moving at a velocity $c_{s}$ along a ring of radius $R(t)$. Accounting for the expansion of the radius $R(t)$ (or a variable velocity $c_{s}$ as detailed in the supplemental material (SM) SuppMat ), we describe such an expanding ring using a Friedmann–Lemaître–Robertson–Walker metric Lachièze-Rey and Luminet (1995); Davies and Fulling (1977):
@@
where $\mathrm{d}x=R_{0}\mathrm{d}\theta$, and the scale factor $a(t)=R(t)/R_{0}$, see Fig. 1. Analogs of such an expanding universe have been proposed when considering the azimuthal mode of a ring of a Bose-Einstein condensate submitted to a forced radial expansion Weinfurtner (2005); Jain et al. (2007a); Eckel et al. (2018); Llorente and Plata (2019).
In a static finite system with $R(t)=R_{0}$ and in the low temperature regime such that $k_{B}T\ll\hbar c_{s}/R$, the thermodynamic properties of a massless relativistic system are controlled by a single energy scale, the Casimir energy density $\varepsilon_{\mathcal{C}}$, set by the size $L=2\pi R$. In particular the energy density and pressure satisfy Francesco et al. (2012)
@@
with $\mathcal{C}_{w}=c_{+}+c_{-}$ where the $c_{\pm}$ are the central charges of the left and right moving excitations of the theory. If the system size or the velocity is slowly varied in time, one expects thermodynamic observables to be set by the instantaneous Casimir energy density $\varepsilon_{\mathcal{C}}(t)$ defined in (2) with $R(t)$. However, a new length scale appears in the problem, the spacetime curvature $\mathcal{R}$. A new energy scale is associated to it, set by the speed of variation of the size $R(t)$:
@@
The emergence of this new characteristic energy scale can be derived from the modification of the vacuum properties of the associated quantum field theory. These modifications are captured by the conformal and gravitational anomalies: once quantized, the massless theory ceases to conserve the vacuum expectation value of the trace and covariant derivative of the stress-energy tensor $\mathcal{T}^{\mu\nu}$ Capper and Duff (1974); Capper et al. (1974); Bertlmann (2000):
@@
with $\mathcal{C}_{g}=c_{+}-c_{-}$. The solutions of equations (Anomalous dynamical Casimir effect in an expanding ring) for a dynamical metric of the form (1) can be decomposed as a sum (see the SM SuppMat ) $\mathcal{T}=\mathcal{T}_{cl}+\mathcal{T}_{q}$. $\mathcal{T}_{cl}$ is the classical momentum energy tensor which satisfies the classical equation of motion (Anomalous dynamical Casimir effect in an expanding ring) with $\mathcal{R}=0$. For an homogeneous initial vacuum state, it satisfies
@@
whose amplitude is set by the new energy scales originating from the gravitational anomalies: $\varepsilon_{\mathcal{R}}$ defined in Eq. (3) and $\varepsilon_{\bar{\mathcal{R}}}(t)=\frac{\hbar c_{s}}{48\pi}\bar{\mathcal{R}}$, with ¯$\bar{\mathcal{R}}=-2c_{s}^{-2}\left(\partial_{t}R/R\right)^{2}$. While $\varepsilon_{\mathcal{R}}$ controls the energy-pressure anisotropy $\varepsilon-p=\mathcal{C}_{w}\varepsilon_{\mathcal{R}}$, both new energy scales set the corrections to the Casimir energy density $\varepsilon+p=\mathcal{C}_{w}(\varepsilon_{\mathcal{C}}-\varepsilon_{\mathcal{R}}+\varepsilon_{\bar{\mathcal{R}}})$. Note that the diagonal part of the tensor (6), related to the conformal anomaly, was identified in Davies and Fulling (1977); Xie et al. (2023). However in these references, the authors focused on the evolution of the energy density in the ring, but did not discuss the behavior of the pressure and the associated instability, which are the core of the physics discussed here.
These corrections to the momentum energy tensor induced by the time variations of the metric appear similar to those obtained in another context when considering inhomogeneous stationary metrics either close to a black hole Christensen and Fulling (1977), or in a wire with an inhomogeneous temperature Bermond et al. (2022). However, in our case, the variation of the metric in time and not in space leads to an change of the corrections to the energy density $\varepsilon$ and to the pressure $p$, and thus to new physical phenomena which we explore in the reminder of this paper.
[…]
From Eqs. (5,6), we get the expression of the energy and momentum density, the pressure, and the energy current of the azimuthal modes as
@@
The instantaneous Casimir energy density $\varepsilon_{\mathcal{C}}(t)$, defined in (2), is therefore modified when either the radial velocity $\partial_{t}R$ is comparable to the particle velocity $c_{s}$, or the acceleration $\partial_{t}^{2}R$ is comparable to $c^{2}_{s}/R$.
The parameters of Fig. 2 are motivated by the experimental work of S. Eckel et al. on a ${}^{23}$a Bose-Einstein condensate described in Eckel et al. (2018). For a slow expansion of the condensate, represented in Fig. 2(a) for $\tau\geq 8$ms, the ratio $\partial_{t}R/c_{s}$ and thus $\varepsilon_{\bar{\mathcal{R}}}$ are negligible. As a consequence, a single anomalous energy scale $\varepsilon_{\mathcal{R}}$ controls the correction to the thermodynamics. The energy density is well approximated by the instantaneous Casimir energy $\varepsilon\approx-\frac{1}{2}\mathcal{C}_{w}\varepsilon_{\mathcal{C}}(t)$ while the pressure is corrected to $p\approx-\varepsilon_{\mathcal{C}}(t)-\mathcal{C}_{w}\varepsilon_{\mathcal{R}}$.
For faster expansions of the condensate both energy scales $\varepsilon_{\mathcal{R}}$ and $\varepsilon_{\bar{\mathcal{R}}}$ are sizable. While a small time dependent correction of the energy density appears as shown in Fig. 2(a), the more striking signature of the anomalous dynamical Casimir effect concerns the evolution of the pressure, represented in Fig. 2(b), which becomes non-monotonous. Such an oscillatory pressure is characteristic of an instability in the system. This instability occurs for negative values of $\partial_{t}p$, which is plotted as a function of time in a rescaled manner in Fig. 2(c) for various $\tau$. The instability arises for fast enough expansions of the ring satisfying $2c_{s}\tau<\sqrt{R_{0}^{2}+R_{1}^{2}+6R_{0}R_{1}}$. In particular, if $c_{s}\tau<\sqrt{2}R_{0}$ they arise for any amplitude $R_{1}/R_{0}$ of the expansion.
The presence of an instability, as well as the difference between energy and pressure, also have consequences on the external work required to drive the expansion of the ring. The work is deduced from the total energy of the vacuum along the ring which reads
The resulting $E(t)$ clearly departs from the instantaneous Casimir energy $E_{\mathcal{C}}(t)=\hbar c_{s}\mathcal{C}_{w}/(24R(t))$, as shown in Fig. 3(a). This departure, which corresponds to the change in the work necessary to drive the radius of the ring, is a signature of the anomalous dynamical Casimir effect discussed in this paper. Accounting for the time-dependence of both energy and volume, the pressure $p=-\mathrm{d}E/\mathrm{d}V=-\partial_{t}E/\partial_{t}R$ can be deduced from (9) and is exactly (7b). The stability condition can be inferred from the compressibility $\beta=-V^{-1}\partial V/\partial p=-R^{-1}\leavevmode\nobreak\ \partial_{t}R/\partial_{t}p$. An instability develops whenever $\partial_{t}p$ vanishes, which occurs for fast enough expansions of the ring as shown in Fig. 3(b). This compressibility is always negative, a consequence of the Casimir effect on a finite size system Cho et al. (2022).
Finally, should we have considered a chiral quantum ring, defined as one where $c_{+}\neq c_{-}$ corresponding to $\mathcal{C}_{g}\neq 0$, the current conservation equation will be modified into Eq. (4b) as a consequence of the gravitational anomaliesBermond et al. (2022). The anomalous dynamical Casimir effect manifests itself as a non-monotonous evolution of the energy current as the size is increased, as inferred from Eq. (7c).
@@
In this work, we have described a new effect, the anomalous dynamical Casimir effect, that alters the groundstate of a massless theory or vacuum in an expanding ring geometry. Because of the anomalous dynamical Casimir effect the required work profile to implement this expansion becomes non-monotonous for fast drives. Such a profile is required to counteract the mechanical instability of the ring manifested by a diverging compressibility during the expansion. The modification of the thermodynamics of the vacuum is, in the language of quantum field theory, due to the conformal and gravitational anomalies. In the SM SuppMat we have also shown that a similar instability occurs then the speed of excitations $c_{s}$ is varied over time, an alternative protocol to varying the ring’s radius as described in the main text. We have shown that the anomalous dynamical Casimir effect is distinct from the dynamical Casimir effect: it originates from the non-conservation equations induced by the conformal and gravitational anomalies (Anomalous dynamical Casimir effect in an expanding ring), in contrast to the momentum energy tensor of the dynamical Casimir effect, which satisfies the standard conservation equations. The possibility to relate this anomalous dynamical Casimir effect to the origin of dark energy along the lines of Mottola (2011) is a fascinating yet open question.
@@
This new phenomenon can be experimentally tested, e.g. in Bose-Einstein condensates, or in quantum circuits realized at the edge of quantum Hall samples. Indeed, while the experiment by Eckel et al Eckel et al. (2018) considered a condensate’s ring submitted to a free expansion and not to a forced one, two experimental manifestations appear in agreement with the instability described in the present paper: first, oscillations of the condensate radius over time are a possible manifestation of the variation of the radiative pressure and the associated instability. Second, the instability can also be related to the development of vortices observed at later times.
Finally, let us mention a few limits of our approach. Our analysis is based on the thermodynamics of the vacuum state. Studying the evolution of the system beyond this adiabatic hypothesis is a stimulating perspective, see e.g. Carusotto et al. (2010). The relation with a Bose-Einstein condensate further requires to neglect the variation of the speed of sound during the evolution. A more accurate description of an experimental situation would require a mixed protocol where both $R(t)$ and $c_{s}(t)$ are driven in time. Additionally, this would also require an extension of our work by incorporating the weak coupling between the 1D azimutal dynamics and the other three-dimensional degrees of freedom.

### v3

@@ Abstract
The Casimir effect is one of the most striking macroscopic manifestations of the quantum nature of vacuum casimir1948attraction. It emerges when this vacuum is subject to boundary conditions, e.g. two metallic plates separated by an electromagnetic vacuum experience an attractive Casimir force bordag2009advances; mostepanenko1997casimir; bressi2002measurement. This effect has been tested experimentally using metals in vacuumbressi2002measurement; PhysRevLett.78.5; PhysRevLett.81.4549 and turned to be repulsive by replacing vacuum by a fluid Munday2009. The Casimir effect generalizes to the case of dielectric plates lifshitz1992theory; Klimchitskaya2009; Woods2016, liquid crystals ajdari1992pseudo, metamaterialsBoyer1974; Kenneth02; Henkel_2005; Leonhardt_2007; Rosa08, liquids and gases PhysRevB.98.201408, and relativistic zhabinskaya2008casimir; shytov2009long; lu2021casimir and topological Grushin2011a; Grushin2011b; Rodriguez-Lopez2014 matter. When the boundary conditions are time-dependent, the increase of vacuum energy and associated properties lead to the so-called dynamical Casimir effects moore1970quantum; dodonov2020fifty; wilson2011observation; jaskula2012acoustic; lahteenmaki2013dynamical; johansson2009dynamical; vezzoli2019optical.
@@
The Casimir effect also affects groundstate properties in a finite geometry without boundaries. For example, when a massless field is confined in a ring of size $L$ in dimension $D=1$, the finite-size energy correction is a manifestation of the Casimir effect francesco2012conformal. In this paper, we consider the fate of this Casimir effect when the size of this ring is rapidly increased in time. Such a system is an analog of a 1D universe in expansion. Analogous expanding universes can be implemented experimentally for example at the edge of a quantum Hall sample with a time-dependent geometry nambu2023analog, and in Bose-Einstein condensate by modulating either their shape fedichev2004cosmological; weinfurtner2005analogue; eckel2018rapidly; llorente2019expanding or their scattering length jain2007analog.
@@
In this work we show that when the expansion of the ring is fast enough the Casimir effect has an additional contribution, distinct from the dynamical Casimir effect resulting from dynamical boundary conditions. Because of this difference, and its link to quantum anomalies bertlmann2000anomalies that we now introduce, we refer to it as the anomalous dynamical Casimir effect. The onset scale of this effect is set by the curvature of spacetime which modifies the ground state properties. In the present case, this curvature of spacetime originates from the time dependence of the spatial curvature of the ring, i.e. its radius. In practice, we account for the modifications of the vacuum properties by this spacetime curvature through the conformal anomaly of the quantum field theory for the energy density and pressure, and the gravitational anomaly for the energy current and momentum bertlmann2000anomalies. We stress that the curved spacetime description of the dynamical Casimir effect moore1970quantum; fulling1976radiation relies on a metric with a vanishing curvature davies1977quantum, unrelated to a conformal anomaly. Hence, the anomalous dynamical Casimir effect, originating from quantum field theory anomalies, is thus distinct from the dynamical Casimir effect in a quantum wire with a varying length or boundary conditions.
@@
In contrast, the anomalous dynamical Casimir effect has a similar origin to that of the Hawking or Unruh radiations. All of these phenomena originate from a curvature of spacetime christensen1977trace; robinson2005relationship; banerjee2008hawking; banerjee2009hawking; gim2015quantal; bermond2022anomalous, originating from either a spatial variation of the metric unruh1981experimental for the Hawking effect, or a dynamical origin for both the Unruh effect and the present anomalous dynamical Casimir effect. The anomalous dynamical Casimir effect opens an alternative to test the effect of a spacetime curvature on the vacuum properties, in addition to the experimental realization of the Hawking effect  unruh1981experimental; garay2000sonic; garay2001sonic; unruh2007quantum; barcelo2001analogue; jain2007quantum; macher2009black; steinhauer2014observation; steinhauer2016observation. While the experimental signatures of these phenomena typically rely on the measurement of the particle density and its correlations jain2007quantum; steinhauer2014observation; steinhauer2016observation; johansson2009dynamical; jaskula2012acoustic, here we show that a remarkable feature of the anomalous dynamical Casimir effect appears as a mechanical instability, related to the behavior of the system’s radiative pressure $p$.
@@
We consider a massless field of excitations moving at a velocity $c_{s}$ along a ring of radius $R(t)$. Accounting for the expansion of the radius $R(t)$ (or a variable velocity $c_{s}$ as detailed in the supplemental material (SM) SuppMat), we describe such an expanding ring using a Friedmann–Lemaître–Robertson–Walker metric LACHIEZEREY1995135; davies1977quantum:
@@
where $\mathrm{d}x=R_{0}\mathrm{d}\theta$, and the scale factor $a(t)=R(t)/R_{0}$, see Fig. 1. Analogs of such an expanding universe have been proposed when considering the azimuthal mode of a ring of a Bose-Einstein condensate submitted to a forced radial expansion weinfurtner2005analogue; jain2007analog; eckel2018rapidly; llorente2019expanding.
In a static finite system with $R(t)=R_{0}$ and in the low temperature regime such that $k_{B}T\ll\hbar c_{s}/R$, the thermodynamic properties of a massless relativistic system are controlled by a single energy scale, the Casimir energy density $\varepsilon_{\mathcal{C}}$, set by the size $L=2\pi R$. In particular the energy density and pressure satisfy francesco2012conformal
@@
with $\mathcal{C}_{w}=c_{+}+c_{-}$ where the $c_{\pm}$ are the central charges of the left and right moving excitations of the theory.
[…]
@@
If the system size or the velocity is slowly varied in time, one expects thermodynamic observables to be set by the instantaneous Casimir energy density $\varepsilon_{\mathcal{C}}(t)$ defined in (2) with $R(t)$. However, a new time-scale characterizing the expansion appears in the problem, $\partial_{t}R/R$. The associated energy-scale is set by the speed of variation of the size $R(t)$:
$\varepsilon^{(a)}(t)=\frac{\hbar}{48\pi c_{s}}\left(\frac{\partial_{t}R}{R}\right)^{2}$ (3)
As a consequence we expect an energy density given by their sum:
$\varepsilon=-\frac{\mathcal{C}_{w}}{2}\left(\varepsilon_{\mathcal{C}}+\varepsilon^{(a)}\right)=-\frac{\mathcal{C}_{w}}{2}\varepsilon_{\mathcal{C}}\left(1+\left(\frac{\partial_{t}R}{c_{s}}\right)^{2}\right)\,.$ (4)
As we will show later, it can also be inferred more rigorously from the modifications of the vacuum properties induced by a non-zero spacetime curvature known as gravitational anomalies capper1974one; capper1974photon; bertlmann2000anomalies. As a consequence, when the radial velocity $\partial_{t}R$ becomes comparable with the longitudinal sound velocity $c_{s}$, we expect the instantaneous Casimir energy density to be modified by the expansion.
Similarly, we expect the radiative pressure to be altered by a fast expansion. Assuming a fast local equilibration, this radiative pressure satisfies
$p=-\frac{\partial_{t}E}{\partial_{t}R}$ (5)
relating the total energy in the system $E=2\pi R\varepsilon$ to the radiative pressure, such that
$p=-\frac{\mathcal{C}_{w}}{2}\varepsilon_{\mathcal{C}}\left(1+\left(\frac{\partial_{t}R}{c_{s}}\right)^{2}-2\frac{R\partial_{t}^{2}R}{c_{s}^{2}}\right)\,.$ (6)
Moreover, we notice that contrarily to the static case, the radiative pressure and the energy density differ: $\varepsilon-p=\varepsilon^{(b)}=-\mathcal{C}_{w}\frac{\hbar}{24\pi c_{s}}\frac{\partial_{t}^{2}R}{R}$. This manifest the breaking of the scale symmetry by the emergent energy scales $\varepsilon^{(a)}$ and $\varepsilon^{(b)}$.
The parameters of Fig. 2 are motivated by the experimental work of S. Eckel et al. on a 23Na Bose-Einstein condensate described in eckel2018rapidly. For a slow expansion of the condensate, represented in Fig. 2(a) for $\tau\geq 8$ms, the ratio $\partial_{t}R/c_{s}$ and thus $\varepsilon^{(a)}$ are negligible. As a consequence, a single anomalous energy scale $\varepsilon^{(b)}$ controls the correction to the thermodynamics. The energy density is well approximated by the instantaneous Casimir energy $\varepsilon\approx-\frac{1}{2}\mathcal{C}_{w}\varepsilon_{\mathcal{C}}(t)$ while the pressure is corrected to $p\approx-\frac{\mathcal{C}_{w}}{2}\varepsilon_{\mathcal{C}}(t)-\mathcal{C}_{w}\varepsilon^{(b)}$.
For faster expansions of the condensate both energy scales $\varepsilon_{\mathcal{R}}$ and $\varepsilon_{\bar{\mathcal{R}}}$ are sizable. While a small time dependent correction of the energy density appears as shown in Fig. 2(a), the more striking signature of the anomalous dynamical Casimir effect concerns the evolution of the pressure, represented in Fig. 2(b), which becomes non-monotonous. Such an oscillatory pressure is characteristic of an instability of the linear modes, which we expect to manifest itself by the excitation of higher energy modes.
The stability condition can be inferred from the compressibility $\beta=-V^{-1}\partial V/\partial p=-R^{-1}\leavevmode\nobreak\ \partial_{t}R/\partial_{t}p$. An instability develops whenever $\partial_{t}p$ vanishes, which occurs for fast enough expansions of the ring as shown in Fig. 3(b) and in a rescaled manner in Fig. 2(c) for various $\tau$. It appears that the instability arises for fast enough expansions of the ring satisfying $2c_{s}\tau<\sqrt{R_{0}^{2}+R_{1}^{2}+6R_{0}R_{1}}$. In particular, if $c_{s}\tau<\sqrt{2}R_{0}$ they arise for any amplitude $R_{1}/R_{0}$ of the expansion. Note that in our case, the compressibility of the linear modes is always negative, a consequence of the Casimir effect on a finite size system cho2022quantum. The presence of an instability, as well as the difference between energy and pressure, also have consequences on the external work required to drive the expansion of the ring. The work is deduced from the total energy of the vacuum along the ring which reads
The resulting $E(t)$ clearly departs from the instantaneous Casimir energy $E_{\mathcal{C}}(t)=\hbar c_{s}\mathcal{C}_{w}/(24R(t))$, as shown in Fig. 3(a). This departure, which corresponds to the change in the work necessary to drive the radius of the ring, is a signature of the anomalous dynamical Casimir effect discussed in this paper.
@@
The emergence of the new energy scales $\varepsilon^{(a)}$ and $\varepsilon^{(b)}$ and their effect on out-of-equilibrium thermodynamics quantities can be derived from the modified vacuum properties of the associated quantum field theory, in the presence of a curved spacetime. These modifications are captured by the conformal and gravitational anomalies: once quantized in a curved spacetime with a scalar curvature $\mathcal{R}\neq 0$, the massless theory ceases to conserve the vacuum expectation value of the trace and covariant derivative of the stress-energy tensor $\mathcal{T}^{\mu\nu}$ capper1974one; capper1974photon; bertlmann2000anomalies:
@@
with $\mathcal{C}_{g}=c_{+}-c_{-}$. The solutions of equations (9) for a dynamical metric of the form (1) can be decomposed as a sum (see the SM SuppMat) $\mathcal{T}=\mathcal{T}_{cl}+\mathcal{T}_{q}$. $\mathcal{T}_{cl}$ is the classical momentum energy tensor which satisfies the classical equation of motion (9) with $\mathcal{R}=0$. For an homogeneous initial vacuum state, it satisfies
@@
whose amplitude is set by the new energy scales originating from the gravitational anomalies
fixed by the spacetime scalar curvature $\mathcal{R}$ and
$\varepsilon_{\bar{\mathcal{R}}}(t)=\frac{\hbar c_{s}}{48\pi}\bar{\mathcal{R}}\,,\quad\bar{\mathcal{R}}=-2c_{s}^{-2}\left(\frac{\partial_{t}R}{R}\right)^{2}\,.$ (13)
fixed by its time average ¯$\bar{\mathcal{R}}=1/R^{2}\int\mathcal{R}\partial_{t}R$. While $\varepsilon_{\mathcal{R}}$ controls the energy-pressure anisotropy $\varepsilon-p=\mathcal{C}_{w}\varepsilon_{\mathcal{R}}$, both new energy scales set the corrections to the Casimir energy density $\varepsilon+p=\mathcal{C}_{w}(\varepsilon_{\mathcal{C}}-\varepsilon_{\mathcal{R}}+\varepsilon_{\bar{\mathcal{R}}})$. Note that the diagonal part of the tensor (11), related to the conformal anomaly, was identified in davies1977quantum; xie2023optomechanical. However in these references, the authors focused on the evolution of the energy density in the ring, but did not discuss the behavior of the pressure and the associated instability, which are the core of the physics discussed here.
These corrections to the momentum energy tensor induced by the time variations of the metric appear similar to those obtained in another context when considering inhomogeneous stationary metrics either close to a black hole christensen1977trace, or in a wire with an inhomogeneous temperature bermond2022anomalous. However, in our case, the variation of the metric in time and not in space leads to an change of the corrections to the energy density $\varepsilon$ and to the pressure $p$, and thus to new physical phenomena which we explore earlier in this paper.
From Eqs. (10,11), we get the expression of the energy and momentum density, the pressure, and the energy current of the azimuthal modes as
@@
The instantaneous Casimir energy density $\varepsilon_{\mathcal{C}}(t)$, defined in (2), is therefore modified when either the radial velocity $\partial_{t}R$ is comparable to the particle velocity $c_{s}$, or the acceleration $\partial_{t}^{2}R$ is comparable to $c^{2}_{s}/R$. A comparison between (12,13) and (3) reveals the identity $\varepsilon_{\mathcal{R}}\equiv\varepsilon^{(b)}$ and $\varepsilon_{\bar{\mathcal{R}}}\equiv\varepsilon^{(a)}$. Finally, had we considered a chiral quantum ring, defined as one where $c_{+}\neq c_{-}$ corresponding to $\mathcal{C}_{g}\neq 0$, the current conservation equation will be modified into Eq. (9b) as a consequence of the gravitational anomalies bermond2022anomalous. In this context, the anomalous dynamical Casimir effect manifests itself as a non-monotonous evolution of the energy current as the size is increased, as inferred from Eq. (14c).
@@
In this work, we have described a new effect, the anomalous dynamical Casimir effect, that alters the groundstate of a massless theory or vacuum in an expanding ring geometry. Because of the anomalous dynamical Casimir effect the required work profile to implement this expansion becomes non-monotonous for fast drives. Such a profile is required to counteract the mechanical instability of the ring manifested by a diverging compressibility during the expansion. The modification of the thermodynamics of the vacuum is, in the language of quantum field theory, due to the conformal and gravitational anomalies. In the SM SuppMat we have also shown that a similar instability occurs then the speed of excitations $c_{s}$ is varied over time, an alternative protocol to varying the ring’s radius as described in the main text. We have shown that the anomalous dynamical Casimir effect is distinct from the dynamical Casimir effect: it originates from the non-conservation equations induced by the conformal and gravitational anomalies (9), in contrast to the momentum energy tensor of the dynamical Casimir effect, which satisfies the standard conservation equations. The possibility to relate this anomalous dynamical Casimir effect to the origin of dark energy along the lines of Mottola_2011, and to the effects of transplanckian physics on the expansion dynamics along the lines of weinfurtner2007trans; cha2017probing are fascinating open questions.
@@
This new phenomenon can be experimentally tested, e.g. in Bose-Einstein condensates, or in quantum circuits realized at the edge of quantum Hall samples. Indeed, while the experiment by Eckel et al eckel2018rapidly considered a condensate’s ring submitted to a free expansion and not to a forced one, two experimental manifestations appear in agreement with the instability described in the present paper: first, oscillations of the condensate radius over time are a possible manifestation of the variation of the radiative pressure and the associated instability. Second, the instability can also be related to the development of vortices observed at later times.
Finally, let us mention a few limits of our approach. Our analysis is based on the thermodynamics of the vacuum state. Studying the evolution of the system beyond this regime is a stimulating perspective, see e.g. carusotto2010density. Moreover, in this work, we described the system using local thermodynamic quantities, following the assumption within the general hydrodynamics framework of a fast local equilibration castro2016emergent; essler2016quench; bernard2016conformal. In other words, we are assuming that the local equilibration time is much smaller than any other time scale of the problem, such as $R/\partial_{t}R$ and $R/c_{s}$.
Relating our results to Bose-Einstein condensates requires extra care. Indeed, hydrodynamic equations in a quantum fluid identify with the equation of motion of quantum fields in a curved spacetime following two crucial assumtions. First, as described in the supplemental materials SuppMat, the dispersion relation can only be approximated as linear up to the healing wavelength of the condensate and second, we need to neglect the time dependence of the speed of sound during the evolution eckel2018rapidly. It is possible to fulfill both assumptions experimentally. First, if we focus on ground state properties of the system, a description based on the linear dispersive low-energy modes appears sufficient unless instabilities excite higher energy modes. Then, while in any experimentally relevant protocol the speed of sound is non-constant due to the change in the background condensed density $n_{0}$ scaling following eckel2018rapidly as $c_{s}\propto R^{-2/7}$, it is possible to counterbalance this effect by modulating the interaction length using e.g. Feshbach resonances vogels1999feshbach; inouye1998observation. Additionally, a full description of the problem would also require an extension of our work by incorporating the weak coupling between the 1D azimutal dynamics and the other three-dimensional degrees of freedom.
Acknowledgements.